OpenStudy (anonymous):

solve equation 8= c/-4+6

OpenStudy (anonymous):

8=c/2 multiply by 2 16=c

mathslover (mathslover):

ok so first of all : \[\huge{\textbf{Welcome to openstudy}}\] is this the question ? \[\huge{8=\frac{c}{-4+6}}\]or \[\huge{8=\frac{c}{-4}+6}\]

OpenStudy (anonymous):

Thanks, its the second question

OpenStudy (anonymous):

2=c/-4 multiply by -4 -8=c

mathslover (mathslover):

ok : \[\huge{8=\frac{c}{-4}+6}\] \[\huge{8=\frac{-c}{4}+6}\] \[\huge{8=\frac{-c+24}{4}}\] \[\huge{8*4=-c+24}\] \[\huge{32=-c+24}\] \[\huge{32-24=-c}\] U can solve further now , i hope

mathslover (mathslover):

just remember :\[\huge{-a=-b}\]then \[\huge{a=b}\]

OpenStudy (anonymous):

thanks!